# Quasinormal modes and bound states of massive scalar fields in generalized Bronnikov-Ellis wormholes

### 23 Feb 2024 v1

@@ I Introduction
Recently there are many efforts to understand modifications of GR which are generated by the presence of a scalar field coupled to gravity. This coupling has important implications in local and global solutions in these theories known as scalar-tensor theories Fujii:2003pa . One of the best studied scalar-tensor theories is provided by the Horndeski Lagrangian Horndeski:1974wa . The Horndeski theory has been studied in short and large distances. In this theory black hole solutions were found Kolyvaris:2011fk ; Rinaldi:2012vy ; Kolyvaris:2013zfa ; Babichev:2013cya ; Charmousis:2014zaa , known as Galilean black holes, and also wormhole geometries were generated Korolev:2014hwa ; Korolev:2020ohi ; Korolev:2020yyy .
@@
An important issue of the compact objects formed by black holes or wormholes is their stability. The relativistic collision of two of these compact objects produces gravitational waves which will help us to understand better the gravitational interaction. However, the recent LIGO detections Abbott:2016blz -TheLIGOScientific:2017qsa do not yet probe the detailed structure of spacetime beyond the photon sphere. Future observations may detect the ringdown phase, which is governed by a series of damped oscillatory modes at early times, which are known as quasinormal modes (QNMs) Vishveshwara:1970zz -Cardoso:2016rao Future Gravitational observations may provide information on different compact objects than black holes, objects without event horizons which are known as exotic compact objects (ECOs) Mazur:2001fv -Holdom:2016nek . Perturbations of black holes and wormholes were extensively studied in the literature. In the case that a test scalar perturbs black hole compact objects the partially reflected waves from the photon sphere (PS) mirror off the AdS boundary and re-perturb the PS to give rise to a damped beating pattern. In the case of a wormhole do not decay with time but have constant and equal amplitude to that of the initial ringdown. Therefore the knowledge of QNMs and quasinormal frequencies (QNFs) can give us important information on the properties of compact objects and distinguish their nature. The QNMs give an infinite discrete spectrum which consists of complex frequencies, $\omega=\omega_{R}-i\omega_{I}$, where the real part $\omega_{R}$ determines the oscillation timescale of the modes, while the complex part $\omega_{I}$ determines their exponential decaying timescale (for a review on QNM modes see Kokkotas:1999bd ; Berti:2009kk ).
@@
In the case that the probe scalar field is massive, a new mass scale is introduced and then a different behaviour was found Konoplya:2004wg ; Konoplya:2006br ; Dolan:2007mj ; Tattersall:2018nve , at least for the overtone $n=0$. If the mass of the scalar field is light, then the longest-lived QNMs are those with a high angular number $\ell$, whereas if the mass of the scalar field is large the longest-lived modes are those with a low angular number $\ell$. One is expecting this behaviour because the fluctuations of the probe massive scalar field can maintain the QNMs to live longer even if the angular number $\ell$ is large, introducing in this way an anomaly of the decaying modes depending on the mass of the scalar field exceeding a critical value or not. Also, if a cosmological constant is introduced an anomalous behaviour of QNMs was found in Aragon:2020tvq because there is interplay of the mass of the scalar field and the value of the cosmological constant. If the background metric is the Reissner-Nordström and the probe scalar field is massless Fontana:2020syy of massive Gonzalez:2022upu depending on critical values of the charge of the black hole, the charge of the scalar field and its mass an anomalous behaviour was also found. The decay modes of QNMs in various setups were studied in Aragon:2020teq -Aragon:2021ogo .
@@
In Gonzalez:2022ote it was shown that if the probe scalar field is massive then an anomalous behaviour of QNMs was also observed if the background metric defines a wormhole geometry. It was shown that in the case of Bronnikov-Ellis wormhole Ellis:1973yv there is a critical mass of the scalar field beyond which the anomalous decay rate of the QNMs is present. In the case of the Morris-Thorne wormhole Morris:1988tu the anomalous decay rate of the QNMs is not present at least for the fundamental modes. In Azad:2022qqn polar perturbations of static Bronnikov-Ellis wormholes were studied.
@@
We also studied the QNFs in a wormhole spacetime supported by a phantom scalar field in which the scalar curvature is generalized to an $f(R)$ function Karakasis:2021lnq ; Karakasis:2021rpn ; Tang:2019qiy ; Karakasis:2021tqx . In this model the scalar field self-interacts with a mass term potential which is derived from the scalar equation and in the resulting $f(R)$ model the scalar curvature is modified by the presence of the scalar field and it is free of ghosts and avoids the tachyonic instability.
@@ IV Generalized Bronnikov-Ellis wormhole
$(2M-r)\left(b_{0}^{2}-r^{2}\right)R^{\prime\prime}(r)+\frac{\left(-3Mb_{0}^{2}+Mr^{2}+b_{0}^{2}r\right)}{r}R^{\prime}(r)+$
$\frac{M\left(\left(2m^{2}r^{2}+2\ell^{2}+2\ell-1\right)r^{2}+3b_{0}^{2}\right)-r\left(r^{2}\left(\ell^{2}+\ell+r^{2}(m^{2}-\omega^{2})\right)+b_{0}^{2}\right)}{r^{2}}R(r)=0\,,$ (21)
@@ IV.1 Quasinormal frequencies
It is convenient to consider the following ansatz for the symmetric solutions of the radial equation (IV) which satisfies the desired boundary conditions
@@
For the fundamental QNFs ($n=0$), we can observe an inverted behaviour of $Im(\omega)b_{0}$, that is, when $mb_{0}>m_{c}b_{0}$, $Im(\omega)b_{0}$ increases with $\ell$, representing an anomalous decay rate of the QNMs. Conversely, for $mb_{0}<m_{c}b_{0}$, $Im(\omega)b_{0}$ decreases with increasing $\ell$ for $n\geq\ell$. The critical value $m_{c}b_{0}$ rises with an increase in the parameter $M/b_{0}$, and the frequency of the oscillation increases with both the angular number and the parameter $mb_{0}$. The frequencies all have a negative imaginary part, which means that the propagation of massive scalar fields is stable in this background. Some numerical values of the QNFs obtained with CFM are in appendix A. Moreover, in appendix B we compare the fundamental QNFs and the first overtone using the CFM and the WKB approximation, and we see a poor agreement with the fundamental QNF but a good agreement for high values of $\ell$ as expected, because it is known that the WKB method only yields accurate results for high values of $\ell$ and for $\ell>n$.
@@ IV.2 Bound States
where $\Omega=\sqrt{m^{2}-\omega^{2}}$. Generally, in a quantum system with a potential well, there are only a few bound states, the number of which depends on the potential parameters. In several cases, there exists only a single bound state. Thus, we just consider the first bound state which is described by the symmetric solution. Therefore, we consider the following ansatz for the solution to the radial equation (IV) which satisfies the desired boundaries conditions
@@ VI Conclusions
This work is partially supported by ANID Chile through FONDECYT Grant Nº 1220871 (P.A.G., and Y. V.).

### 05 Mar 2024 v2

@@ I Introduction
Recently there are many efforts to understand modifications of GR which are generated by the presence of a scalar field coupled to gravity. This coupling has important implications in local and global solutions in these theories known as scalar-tensor theories Fujii:2003pa . One of the best studied scalar-tensor theories is provided by the Horndeski Lagrangian Horndeski:1974wa . The Horndeski theory has been studied in short and large distances. In this theory black hole solutions were found Kolyvaris:2011fk ; Rinaldi:2012vy ; Kolyvaris:2013zfa ; Babichev:2013cya ; Charmousis:2014zaa , known as Galilean black holes, and also wormhole geometries were generated Korolev:2014hwa -Bakopoulos:2023tso
@@
An important issue of the compact objects formed by black holes or wormholes is their stability. The relativistic collision of two of these compact objects produces gravitational waves which will help us to understand better the gravitational interaction. However, the recent LIGO detections Abbott:2016blz -TheLIGOScientific:2017qsa do not yet probe the detailed structure of spacetime beyond the photon sphere. Future observations may detect the ringdown phase, which is governed by a series of damped oscillatory modes at early times, which are known as quasinormal modes (QNMs) Vishveshwara:1970zz -Cardoso:2016rao Future Gravitational observations may provide information on different compact objects than black holes, objects without event horizons which are known as exotic compact objects (ECOs) Mazur:2001fv -Holdom:2016nek . Perturbations of black holes and wormholes were extensively studied in the literature. In the case that a test scalar perturbs black hole compact objects the partially reflected waves from the photon sphere (PS) mirror off the AdS boundary and re-perturb the PS to give rise to a damped beating pattern. In the case of a wormhole do not decay with time but have constant and equal amplitude to that of the initial ringdown. Therefore the knowledge of QNMs and quasinormal frequencies (QNFs) can give us important information on the properties of compact objects and distinguish their nature. The QNMs give an infinite discrete spectrum which consists of complex frequencies, $\omega=\omega_{R}+i\omega_{I}$, where the real part $\omega_{R}$ determines the oscillation timescale of the modes, while the complex part $\omega_{I}$ determines their exponential decaying timescale (for a review on QNM modes see Kokkotas:1999bd ; Berti:2009kk ).
@@
In the case that the probe scalar field is massive, a new mass scale is introduced and then a different behaviour was found Konoplya:2004wg ; Konoplya:2006br ; Dolan:2007mj ; Tattersall:2018nve , at least for the overtone $n=0$. If the mass of the scalar field is light, then the longest-lived QNMs are those with a high angular number $\ell$, whereas if the mass of the scalar field is large the longest-lived modes are those with a low angular number $\ell$. One is expecting this behaviour because the fluctuations of the probe massive scalar field can maintain the QNMs to live longer even if the angular number $\ell$ is large, introducing in this way an anomaly of the decaying modes depending on the mass of the scalar field exceeding a critical value or not. Also, if a cosmological constant is introduced an anomalous behaviour of QNMs was found in Aragon:2020tvq because there is interplay of the mass of the scalar field and the value of the cosmological constant. If the background metric is the Reissner-Nordström and the probe scalar field is massless Fontana:2020syy of massive Gonzalez:2022upu depending on critical values of the charge of the black hole, the charge of the scalar field and its mass an anomalous behaviour was also found. The decay modes of QNMs in various setups were studied in Aragon:2020teq -Becar:2023zbl .
@@
In Gonzalez:2022ote it was shown that if the probe scalar field is massive then an anomalous behaviour of QNMs was also observed if the background metric defines a wormhole geometry. It was shown that in the case of Bronnikov-Ellis wormhole Ellis:1973yv there is a critical mass of the scalar field beyond which the anomalous decay rate of the QNMs is present. In the case of the Morris-Thorne wormhole Morris:1988tu the anomalous decay rate of the QNMs is not present at least for the fundamental modes. In Azad:2022qqn polar perturbations of static Bronnikov-Ellis wormholes were studied and the QNMs of rapidly rotating Bronnikov-Ellis wormholes were recently analyzed in Khoo:2024yeh .
@@
We also studied the QNFs in a wormhole spacetime supported by a phantom scalar field in which the scalar curvature is generalized to an $f(R)$ function Karakasis:2021lnq ; Karakasis:2021rpn ; Tang:2019qiy ; DeFalco:2021ksd ; Karakasis:2021tqx . In this model the scalar field self-interacts with a mass term potential which is derived from the scalar equation and in the resulting $f(R)$ model the scalar curvature is modified by the presence of the scalar field and it is free of ghosts and avoids the tachyonic instability.
@@ IV Generalized Bronnikov-Ellis wormhole
$(2M-r)\left(b_{0}^{2}-r^{2}\right)R^{\prime\prime}(r)+\frac{\left(-3Mb_{0}^{2}+Mr^{2}+b_{0}^{2}r\right)}{r}R^{\prime}(r)+$ (21)
$\frac{M\left(\left(2m^{2}r^{2}+2\ell^{2}+2\ell-1\right)r^{2}+3b_{0}^{2}\right)-r\left(r^{2}\left(\ell^{2}+\ell+r^{2}(m^{2}-\omega^{2})\right)+b_{0}^{2}\right)}{r^{2}}R(r)=0\,,$
@@ IV.1 Quasinormal frequencies
It is convenient to consider the following ansatz for the symmetric solutions of the radial equation (21) which satisfies the desired boundary conditions
@@
For the fundamental QNFs ($n=0$), we can observe an inverted behaviour of $Im(\omega)b_{0}$, that is, when $mb_{0}>m_{c}b_{0}$, $Im(\omega)b_{0}$ increases with $\ell$, representing an anomalous decay rate of the QNMs. Conversely, for $mb_{0}<m_{c}b_{0}$, $Im(\omega)b_{0}$ decreases with increasing $\ell$ for $n\geq\ell$. The critical value $m_{c}b_{0}$ rises with an increase in the parameter $M/b_{0}$, and the frequency of the oscillation increases with both the angular number and the parameter $mb_{0}$. The frequencies all have a negative imaginary part, which means that the propagation of massive scalar fields is stable in this background. Some numerical values of the QNFs obtained with CFM are in appendix A. Moreover, in appendix B we compare the fundamental QNFs and the first overtone using the CFM and the WKB approximation, and we see a good agreement for high values of $\ell$ as expected, because it is known that the WKB method only yields accurate results for high values of $\ell$ and for $\ell>n$.
@@ IV.2 Bound States
where $\Omega=\sqrt{m^{2}-\omega^{2}}$. Generally, in a quantum system with a potential well, there are only a few bound states, the number of which depends on the potential parameters. In several cases, there exists only a single bound state. Thus, we just consider the first bound state which is described by the symmetric solution. Therefore, we consider the following ansatz for the solution to the radial equation (21) which satisfies the desired boundaries conditions
@@ VI Conclusions
This work is partially supported by ANID Chile through FONDECYT Grant Nº 1220871 (P.A.G., and Y. V.). P. A. G. acknowledges the hospitality of the Universidad de La Serena where part of this work was undertaken.